the equation 2x-3x^1/2 + 1 = 0 has two solutions. The sum of the solutions is:
a 1 b 1/2 c 5/4 d 4 e 3/2

Answer by texttutoring(324)

Recall that x^1/2 is the same as Sqrt(x).

Factor:
2x -3x^(1/2) +1 =0
(2x^(1/2) - 1)(x^(1/2) - 1) = 0

So either:

2x^(1/2) - 1 =0
2x^(1/2)=1
x=1/4

OR

x^(1/2) - 1 = 0
x^(1/2) = 1
x= 1

Add 1+1/4 to get 5/4.

The sum of the answers is 5/4.